# Fixture calibration with guaranteed bounds from a few correspondence-free surface points

### 28 Feb 2024 v1

@@ Abstract
Calibration of fixtures in robotic work cells is essential but also time consuming and error-prone, and poor calibration can easily lead to wasted debugging time in downstream tasks. Contact-based calibration methods let the user measure points on the fixture’s surface with a tool tip attached to the robot’s end effector. Most methods require the user to manually annotate correspondences on the CAD model, however, this is error-prone and a cumbersome user experience. We propose a correspondence-free alternative: The user simply measures a few points from the fixture’s surface, and our method provides a tight superset of the poses which could explain the measured points. This naturally detects ambiguities related to symmetry and uninformative points and conveys this uncertainty to the user. Perhaps more importantly, it provides guaranteed bounds on the pose. The computation of such bounds is made tractable by the use of a hierarchical grid on SE(3). Our method is evaluated both in simulation and on a real collaborative robot, showing great potential for easier and less error-prone fixture calibration.
@@ I Introduction
Figure 1: A user moves a robot to measure a few points on a fixture using a calibrated tool-tip attached to the end-effector of the robot. The tool consist of a 3D-printed mount with a metal ball bearing at the tip. The fixture is a 3D-printed MATLAB logo, fixated to the table using additional fixtures.
@@
Existing contact-based methods rely on correspondences to infer the pose. Most methods require the user to manually annotate these correspondences, either in form of point-to-point, point-to-line and/or point-to-plane correspondences. Repeatedly going back and forth between demonstrating points with the robot and annotating correspondences in a computer program is not only a cumbersome user experience; It also requires the user to identify features on the fixture which the user can measure with low uncertainty. This process results in correspondences with uncertainties that are primarily related to human ability, and it is therefore difficult to assume a maximum correspondence error, especially for non-expert users.
@@ IV-A Simulation
When the superset is uni-modal, the bounds and confidence intervals, which can be obtained, become of interest. To this end, we simulate experiments on the MATLAB logo, shown in Figure 1, with varying sample error bounds. The results are shown in Figure 2. The positional errors of the expected poses are lower than the sample error bounds and well below both the confidence intervals and pose bounds. We observe similar results for other fixtures.
@@ IV-B1 Tool design
We design a low-cost point measuring tool which can be seen in Figure 1 and 4. The tool consists of an 3D-printed cone and an off-the-shelf metal ball bearing. The base of the cone is attached to the end-effector of the robot, and an indent at the tip of the cone allows the ball bearing to be glued firmly onto the cone. This way, we obtain both the flexibility of 3D-printing, which makes it easy to incorporate a measuring tip in tool designs, and the tip-tolerances and wear-resistance of a metal ball bearing.
Figure 4: The tool-tip with the ball bearing is shown in black. The fixture is shown in red, and the Minkowski sum of the ball bearing and the fixture is illustrated by a dashed, red line. The center of the ball bearing lies on the Minkowski sum surface independently of the angle of sampling.
@@ IV-B2 Tip calibration
This calibration provides the center of the ball bearing, and it turns out that this is desirable. When sampling points on the fixture surface, it is not clear which point on the ball bearing is in contact with the surface, however, it is clear, that the center of the ball bearing lies on the Minkowski sum [31] of the fixture and the ball bearing. See Figure 4. Since the ball bearing and fixture geometries are known, we can compute this Minkowski sum. We use the CGAL implementation [32]. We set the sample error bound above the largest residual observed in Eq. 17: $b_{\text{s}}{}=1.0~{}\text{mm}$.

### 04 Mar 2024 v2

@@ Abstract
Calibration of fixtures in robotic work cells is essential but also time consuming and error-prone, and poor calibration can easily lead to wasted debugging time in downstream tasks. Contact-based calibration methods let the user measure points on the fixture’s surface with a tool tip attached to the robot’s end effector. Most such methods require the user to manually annotate correspondences on the CAD model, however, this is error-prone and a cumbersome user experience. We propose a correspondence-free alternative: The user simply measures a few points from the fixture’s surface, and our method provides a tight superset of the poses which could explain the measured points. This naturally detects ambiguities related to symmetry and uninformative points and conveys this uncertainty to the user. Perhaps more importantly, it provides guaranteed bounds on the pose. The computation of such bounds is made tractable by the use of a hierarchical grid on SE(3). Our method is evaluated both in simulation and on a real collaborative robot, showing great potential for easier and less error-prone fixture calibration. sites.google.com/view/ttpose
@@ I Introduction
Figure 1: A user moves a robot to measure a few points on a fixture using a calibrated tool-tip attached to the end-effector of the robot. The tool consist of a 3D-printed mount with a 3 mm steel ball at the tip. The fixture is a 3D-printed MATLAB logo, fixated to the table using additional fixtures.
@@
Existing contact-based methods rely on correspondences to infer the pose. Most methods require the user to manually annotate these correspondences, either in form of point-to-point, point-to-line and/or point-to-plane correspondences. Repeatedly going back and forth between demonstrating points with the robot and annotating correspondences in a computer program is not only a cumbersome user experience; It also requires the user to identify features on the fixture which the user can measure with low uncertainty. This process results in uncertainties that are primarily related to human ability, and it is thus difficult to assume a maximum correspondence error, especially for non-expert users.
@@ IV-A Simulation
When the superset is uni-modal, the bounds and confidence intervals which can be obtained, become of interest. To this end, we simulate experiments on the MATLAB logo, shown in Figure 1, with varying sample error bounds. The results are shown in Figure 2. The positional errors of the expected poses are lower than the sample error bounds and well below both the confidence intervals and pose bounds. We observe similar results for other fixtures.
@@ IV-B1 Tool design
We design a low-cost point measuring tool which can be seen in Figure 1 and 4. The tool consists of an 3D-printed cone and an off-the-shelf 3 mm steel ball. The cone is attached to the robot’s end effector, and an indent at the tip of the cone allows the steel ball to be glued firmly onto the cone. This way, we get the flexibility of 3D-printing, for easy incorporation in tool designs, and the tolerances and wear-resistance of a steel ball.
Figure 4: The tool-tip with the steel ball is shown in black. The fixture is shown in red, and the Minkowski sum of the steel ball and the fixture is illustrated by a dashed, red line. The center of the steel ball lies on the Minkowski sum surface independently of the angle of sampling.
@@ IV-B2 Tip calibration
This calibration provides the center of the steel ball, and it turns out that this is desirable. When sampling points on the fixture surface, it is not clear which point on the steel ball is in contact with the surface, however, it is clear, that the center of the steel ball lies on the Minkowski sum [31] of the fixture and the steel ball. See Figure 4. Since the steel ball and fixture geometries are known, we can compute this Minkowski sum. We use the CGAL implementation [32]. We set the sample error bound above the largest residual observed in Eq. 17: $b_{\text{s}}{}=1.0~{}\text{mm}$.